# U2A: Unified Unimodal Adaptation for Robust and Efficient Multimodal Learning

### v1

@@ 5 Discussion and Future Directions
Limitations and future directions. Our approach demonstrates better performance in both complete and missing modality scenarios. But there are still some limitations and open challenges. First, our approach can partially recover the performance drop due to missing modalities. This indicates scope for more effective compensation methods. Second, We observe that certain classes are heavily dependent on a specific modality. This leads to significant performance drop in those classes when that specific modality is missing. It brings up an important question: how to design a better training strategy that encourages more uniform reliance on different modalities while maintaining overall performance in both complete and missing modality scenarios? Finally, future work can also explore developing a general framework for quantifying modality contribution and establishing theoretical bounds for performance drop under different missing modality scenarios.
Figure 5: Per class performance analysis on MM-IMDb dataset.
@@ 6 Conclusion
In this paper, we present a simple yet effective approach for utilizing pre-trained unimodal encoders on multimodal tasks through unified adaptation on multimodal data. Our method reduces the complexity of traditional training strategies and minimizes the number of learnable parameters, while still achieving strong performance. It can accommodate both complete and missing modalities during training and testing. By using a single mask token per modality to estimate missing modality features, we avoid the need for specialized feature generation modules or prompt-tuning techniques. Our extensive evaluations across diverse tasks and datasets demonstrate the robustness and adaptability of our approach in both complete and missing modality scenarios. Our analysis also reveals that different classes depend on specific modalities, providing insights into how the absence of a modality impacts performance across different classes. Overall, our approach provides a robust and efficient solution for multimodal learning that maintains high performance in different real-world scenarios.
@@ 7 Acknowledgment
1 Detailed Dataset Description
@@
Table 6: Performance comparison on MM-IMDb and UPMC Food-101 dataset when modality gets missing during both training and testing. We train and test our model with different modality ratios and report the mean and standard deviation (SD) of three independent runs with different seeds. We report F1-Macro score for MM-IMDb dataset and % accuracy for UPMC Food-101 dataset.
Datasets Mising Rate Train Modality Test Modality ViLT MAP MSP Kim et al. Ours
Image Text Image Text (Mean $\pm$ SD)
MM-IMDb 70% 30% 100% 30% 100% 37.63 47.18 47.45 56.03 60.21 $\pm$ 0.40
@@
2 Implementation Details
@@
3 Different Modality Ratios During Training and Testing
@@
To further evaluate the missing modality robustness of our method, we follow the experimental setup from [25, 48] and train and test our method with different modality ratios. As described in Section 4.4.1, let $\eta$ denote the missing rate. We consider three scenarios: missing-text, missing-image, and missing-both. For the first two cases, $\eta\%$ of the samples contain only images or only texts, while $(1-\eta)\%$ contain image-text pairs. For the missing-both case, we have $\frac{\eta}{2}\%$ text-only, $\frac{\eta}{2}\%$ image-only, and $(1-\eta)\%$ complete pairs. We set $\eta=70$ throughout our experiments following [25, 48].
We compare our method with the following approaches: (1) baseline multimodal model: ViLT [8]; and (2) prompt-based methods: missing-aware prompts (MAP) [24], modality-specific prompts (MSP) [25], and method proposed by Kim and Kim [48]. Other methods do not report performance for this specific setup, and thus we are unable to compare with them.
@@
Table 6 presents the results for different missing modality scenarios during training and testing. To ensure robust evaluation, we report the mean and standard deviation from three independent runs with different seeds. While the method proposed by Kim and Kim [48] performs well in some scenarios, our approach consistently outperforms existing baseline methods on both the MM-IMDb and UPMC Food-101 datasets in most of the cases. These results support our claim that mask tokens effectively estimate missing modality features, enhancing robustness to different missing modality scenarios.
@@
Figure 6: Per class performance analysis on UPMC Food-101 dataset.
4 Analysis of Modality Dependence.
@@
We extend our analysis of the impact of missing modalities on UPMC Food-101 dataset. We train our model on modality-complete data and evaluate it on both modality-complete and modality-incomplete test samples. As shown in Figure 6, we observe that multimodal (image-text) data consistently outperforms unimodal performance across all classes. However, the dependence on specific modalities varies significantly across different food categories.
@@
Classes such as “dumplings”, “Frozen_yogurt”, and “Ice cream” demonstrate a strong dependence on the text modality. These classes perform well with text-only inputs, and performance drops drastically when text modality is missing. In contrast, categories like “tuna_tartare”, “beef_carpaccio”, and “beet_salad” depend heavily on the image modality. Their performance drops significantly when image modality is missing. This observation suggests that some classes rely more on text, while others depend heavily on images for accurate classification.
@@
Interestingly, some classes such as “beef_tartare”, “chicken_quesadilla”, and “croque_madame” exhibit relatively poor performance with unimodal inputs. Performance on these classes improve significantly when we utilize both image and text. Overall, our findings on UPMC Food-101 dataset supports our previous observation, as discussed in Section 4.5, that while multimodal information universally enhances performance, the degree of modality dependence remains heterogeneous across classes.

### v2

@@ 5 Limitations and Future Directions
Below we list some limitations and potential future directions. First, while our approach shows improved performance with missing modalities over existing methods, missing modalities still cause performance drop. Second, we observe that certain classes are heavily dependent on a specific modality, which leads to significant performance drop in those classes when that specific modality is missing. A natural question arises as how can we design a better training strategy that encourages more uniform reliance on different modalities while maintaining overall performance in both complete and missing modality scenarios? Future work can also explore developing a general framework for explaining modality contribution and expected performance drop under different missing modality scenarios.
@@ 6 Conclusion
We presented a simple yet effective framework for utilizing pre-trained unimodal encoders for robust multimodal learning. Our proposed framework enables training of multimodal systems that offer robust performance under different scenarios of missing modalities during training and inference. To compensate for missing modalities, we propose to learn cross-modal proxy tokens (CMPTs) to approximate the missing modality class tokens from the available modality using an alignment loss during training. Instead of training the unimodal encoders from scratch, we only learn rank-1 adapters along with task-specific classifier heads. Our extensive evaluations across diverse tasks and datasets demonstrate the robustness and adaptability of our approach in both complete and missing modality scenarios. Our analysis also reveals that CMPTs can generalize well to different missing scenarios and enhance performance across most of the classes to provide and overall performance boost. Overall, our approach provides a robust and efficient solution for multimodal learning that maintains high performance in different real-world scenarios.
@@ 7 Acknowledgment
S1 Related Work on Multimodal Learning
In this section, we provide a general overview of multimodal learning, which involves integrating information from different input sources to improve performance on downstream tasks. We can broadly categorize multimodal learning approaches into two main categories: designing specialized models and adapting existing models.
Designing specialized models to effectively fuse multiple input sources is a common strategy in multimodal learning. For vision-language tasks, models such as ViLT [21], ViLBERT [27], and Visual-BERT [23] process image patches and text tokens directly within a transformer architecture for tasks like image captioning and visual question answering. MMBT [19] integrates both text and images within a BERT-based framework [7] to model cross-modal dependencies. MBT [33] leverages fusion bottlenecks for efficient modality-specific interaction, particularly in audio-video tasks. Specialized multimodal models have also been proposed for other applications, including action recognition [51, 56], segmentation [53, 49], and sentiment analysis [51, 43]. While these models excel in specific tasks, extending them to new modalities or domains often proves challenging and resource-intensive, as it requires retraining or significant modification of the model architecture.
Adapting existing models offers an alternative approach to multimodal learning, focusing on the use of pre-trained unimodal models and adapting them for multimodal tasks. One such strategy is gradient modulation, where techniques like G-Blend [45] modulate gradients to combine unimodal and multimodal knowledge. OGM-GE [36] improves upon this by emphasizing modality-specific information, enhancing the model’s ability to leverage unique features from each input source. PMR [11] further refines this approach by modulating parameter groups based on the relevance of each modality, while QMF [54] introduces quadratic modulation for adaptive fusion. More recently, MLA [55] proposed alternating adaptation, which toggles between unimodal adaptation and gradient modulation to enable more efficient cross-modal learning. Another notable approach, MMLoRA [9], uses a two-phase strategy: unimodal fine-tuning followed by multimodal adaptation, which helps achieve optimal performance across a variety of tasks. However, while these methods allow for effective cross-modal learning, they often require significant computational resources and learning a large number of parameters to achieve optimal performance on downstream tasks.
@@
Table 6: Performance comparison when modality gets missing during both training and inference. We train and evaluate our model with different modality ratio following [16, 20] and report the mean and standard deviation (SD) of three independent runs with different seeds.
Datasets Missing Rate ($\eta$) Train Inference ViLT MAP MSP Kim et al. CMPT (Ours)
Image Text Image Text [21] [22] [16] [20] (Mean $\pm$ SD)
MM-IMDb (F1-Macro) 70% 30% 100% 30% 100% 37.63 47.18 47.45 56.03 60.21 $\pm$ 0.40
@@
S2 Detailed Dataset Description
(a) Available: Text, Missing: Image.
(b) Available: Image, Missing: Text.
Figure 6: Effectiveness of CMPT in handling missing modalities on UPMC Food-101 dataset. CMPT significantly improves performance when modalities are severely missing. Models are trained on 100% image + 100% text and tested with varying levels of missing modalities.
S3 Implementation Details
@@
S4 Different Modality Ratios During Training and Inference
@@
A robust multimodal learning method should maintain strong performance even when the missing modality ratio during inference differs from that during training. To evaluate the robustness of our approach, we follow the experimental setup of [16, 20] and train and test our model with varying modality ratios. Let $\eta$ denote the missing rate. We consider three scenarios: missing-text, missing-image, and missing-both. In the first two cases, $\eta\%$ of the samples contain only images or only text, while $(1-\eta)\%$ retain both modalities. For the missing-both case, $\frac{\eta}{2}\%$ of the samples contain only texts, $\frac{\eta}{2}\%$ contain only images, and $(1-\eta)\%$ include both modalities. We set $\eta=70$ throughout our experiments, following [16, 20].
@@
As summarized in Table 6, our method consistently outperforms existing state-of-the-art approaches in most scenarios. While the method proposed by Kim et al. [20] performs well in certain cases, our approach demonstrates superior performance across both the MM-IMDb and UPMC Food-101 datasets. To ensure a rigorous evaluation, we report the mean and standard deviation (SD) over three independent runs with different seeds. The low SD values further indicate the stability and reliability of our method. These results substantiate our claim that CMPTs can effectively estimate missing modality features, enhancing robustness to varying missing modality conditions. Other methods reported in Table 1 do not report results for this specific experimental setup, and thus we are unable to compare with them.
S5 Effectiveness of CMPT
We extend the study from Section 4.6.1 to evaluate the effectiveness of CMPT in handling missing modalities on the UPMC Food-101 dataset. We compare it against two baselines: (i) a standard model trained without modality dropout or CMPT (baseline) and (ii) a model trained with modality dropout, where modalities are randomly removed during training to improve robustness to missing data.
As illustrated in Figure 6, the baseline model suffers a sharp performance decline as the proportion of missing modalities increases. While modality dropout improves robustness, it fails to fully recover the lost performance. Incorporating CMPT further enhances performance, achieving a 5.07% improvement over the baseline and 3.81% over modality dropout when the image modality is entirely missing as shown in Figure 6(a). A similar trend is observed when text is missing as shown in Figure 6(b)), where CMPT achieves 25.59% and 11.96% higher accuracy compared to the baseline and modality dropout, respectively. Consistent with our previous findings, CMPT becomes increasingly effective as the proportion of missing data increases, demonstrating its strong capability to mitigate performance degradation in the presence of severe modality loss.
Figure 7: CMPT improves performance across most of the classes when text modality is missing on UPMC Food-101 dataset. However, it fails to accurately predict certain classes (e.g., bibimbap/sashimi), likely due to the image modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
Figure 8: CMPT improves performance across most of the classes when image modality is missing on UPMC Food-101 dataset. However, it fails to accurately predict certain classes (e.g., lobster_roll_sandwitch/cheese_plate), likely due to the text modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
Figure 9: Per class performance analysis on MM-IMDb dataset. Certain classes show strong dependence on a specific modality.
@@
Figure 10: Per class performance analysis on UPMC Food-101 dataset. Certain classes show strong dependence on a specific modality.
S6 Does CMPT Benefit All Classes?
To further evaluate CMPT’s effectiveness, we conduct a per-class performance analysis on the UPMC Food-101 dataset, following the approach from Section 4.6.2.
As shown in Figure 7 and 8, CMPT improves or maintains performance across most of the classes in both text-missing and image-missing scenarios. Notably, it mitigates performance degradation in modality-sensitive classes. For instance, when text is missing, classes such as guacamole, pancakes, and foie_gras experience severe drops in performance, but CMPT effectively recovers much of the loss as shown in Figure 6(b). Similarly, in image-missing scenarios, CMPT provides substantial gains for sushi, cup_cakes, and pizza as shown in Figure 4(a). However, a slight performance drop is observed in certain classes when either text or image is absent, likely due to insufficient information in the available modality alone for those classes. Overall, these results reinforce our previous findings, demonstrating that CMPT enhances performance across most classes, leading to a strong overall performance improvement.
S7 Analysis of Modality Dependence
@@
We observe that the impact of missing modalities varies significantly across classes. Certain classes show strong dependence on a specific modality. In this experiment, we train our model on modality-complete data and evaluate it on both modality-complete and modality-incomplete test samples.
Figure 9 presents per-class performance on the MM-IMDb dataset, revealing heterogeneous effects of missing modalities. For instance, the Drama and Thriller classes maintain consistent performance regardless of missing modalities. In contrast, the Biography and Documentary classes experience significant decline in performance when text is missing, even if images are available. Interestingly, certain classes like Animation show slightly higher performance with single-modality inputs, particularly in image-only scenario. Similar patterns emerge in Sport and Music categories, where text-only samples performs slightly better than complete modality. This discrepancy may stem from the simple sum-fusion strategy we use, which might not effectively model complex modality interactions. These findings indicate that the effect of missing modalities is not uniform across classes, with some being more modality-dependent than others.
Extending this analysis to UPMC Food-101, we observe that multimodal (image-text) data consistently outperforms unimodal inputs across all classes as shown in Figure 10. However, the reliance on specific modalities varies widely among food categories.
@@
Classes such as dumplings, Frozen_yogurt, and Ice cream depend heavily on text, showing strong performance with text-only inputs but experiencing significant drops when text is missing. Conversely, categories like tuna_tartare, beef_carpaccio, and beet_salad rely more on image information, with performance deteriorating notably when images are absent. This observation suggests that some classes rely more on texts, while others depend heavily on images for accurate classification.
@@
Interestingly, some classes such as beef_tartare, chicken_quesadilla, and croque_madame exhibit poor performance with unimodal inputs but improve significantly with both image and text, reinforcing the importance of multimodal data. Performance on these classes improve significantly when we utilize both image and text. Overall, these findings confirm that while multimodal information universally enhances performance, the degree of modality dependence remains heterogeneous across classes.